Can someone please help with this problem I do not understand?
f(x)=x=7 what is f(-15)=

Found 2 solutions by Alan3354, Edwin McCravy:
Answer by Alan3354(69443)

f(x)=x=7
Too many equal signs

Answer by Edwin McCravy(20054)

f(x)=x+7 what is f(-15)=
------------------------


I will assume you made a typo and failed to hold down the shift key,
and typed an = instead of a +.

Start with 

f(x) = x + 7

And everywhere you see x, instead of x write -15 instead of x

and you have

f(-15) = -15 + 7

then you simplify the right side:

f(-15) = -8

and you are done.  

[Note that f(x) does NOT mean fx nor does f(-15) mean -15f.  You leave
f(-15) just as it is and the final answer is f(-15) = -8]  


Edwin